### v1

@@ Abstract
We present an exact static spherisymmetric solution for the Brans-Dicke action sourced by a self-gravitating massless Klein-Gordon helicity–0 field. In contrast to the Maxwell electromagnetic field, a Klein-Gordon field possesses an energy-momentum tensor with non-vanishing trace. Upon a Weyl mapping into the Einstein frame, the transformed Brans-Dicke scalar field takes on the role of a “dilaton” coupled with the Klein-Gordon field. Despite this dilatonic coupling, the field equations of the resulting Einstein-Klein-Gordon-dilaton action are fully soluble when employing the harmonic radial coordinate. The exact solution derived herein can serve as a prototype for future Brans-Dicke gravity studies involving trace-carrying matter fields. Notably, in the limit of $\omega\rightarrow\infty$, the Brans-Dicke scalar field exhibits an anomalous behavior of $\mathcal{O}(1/\sqrt{\omega})$ as opposed to $\mathcal{O}\left(1/\omega\right)$. As a consequence, the solution converges to a spacetime configuration of General Relativity sourced by the original Klein-Gordon field and a free scalar field, the latter of which is the ${\cal O}\left(1/\sqrt{\omega}\right)$ “remnant” of the Brans-Dicke scalar field. Furthermore, we provide a formal mathematical proof substantiating these two conclusions. Although the $\mathcal{O}\left(1/\sqrt{\omega}\right)$ anomaly has been previously discovered for Brans-Dicke vacuum and Brans-Dicke-Maxwell electrovacuum, our findings establish its prevalence in Brans-Dicke gravity regardless of the trace of the energy-momentum tensor of the source. Taken together, the ${\cal O}\left(1/\sqrt{\omega}\right)$ anomaly challenges the conventional belief in the $\mathcal{O}(1/\omega)$ signature commonly associated with Brans-Dicke gravity. In particular, it may have implications in improving the relativistic corrections to Newtonian gravity beyond the weak-field parametrized post-Newtonian formalism.
@@ I Motivation
This scenario is precisely what happens in BD gravity, revealing a non-convergence behavior that has been extensively explored and documented since the 1990s. Specifically, various authors have established that the infinite $\omega$ limit fails to transition the BD vacuum to a vacuum of GR [9, 10, 11, 12, 13, 14, 15, 16, 18, 19, 22, 20, 17, 21, 23]. Instead, it approaches a GR solution sourced by a free massless scalar field; (note: for static spherisymmetric setup, the limiting vacuum is the Fisher-Janis-Newman-Winicour (FJNW) solution  [^1] The solution has been re-discovered several times, and is also known as the Fisher-Bergmann-Leipnik-Janis-Newman-Winicour-Buchdahl-Wyman (FBLJNWBW) solution [23].). Moreover, the convergence rate in the limit is determined to be ${\cal O}\left(1/\sqrt{\omega}\right)$ instead of ${\cal O}\left(1/\omega\right)$. That is to say, the BD scalar field behaves as
@@
with the subleading field $\stackrel{{\scriptstyle 1}}{{\varphi}}$ being non-constant. We shall aptly call $\stackrel{{\scriptstyle 1}}{{\varphi}}$ a “remnant” BD field in the infinite $\omega$ limit.
These findings are two faces of the same coin. Due to the $\stackrel{{\scriptstyle 1}}{{\varphi}}$ associated with the $1/\sqrt{\omega}$ order, the “kinetic” term remains finite, viz.
$\omega\left(\nabla\phi/\phi\right)^{2}\simeq\bigl{(}\ \stackrel{{\scriptstyle 1}}{{\varphi}}\>\bigr{)}^{2}$ (3)
@@
Secondly, we shall reinforce this proof by examining a concrete case: the BD theory sourced by a massless Klein-Gordon helicity-0 field. For the static spherisymmetric setup, this theory proves to be soluble, and the $\mathcal{O}\left(1/\sqrt{\omega}\right)$ signature manifests in the final analytical solution. As a by-product, the exact solution obtained herein can also serve as a prototype for future studies involving the BD action coupled with trace-carrying matter fields.
@@
Our paper is organized as follows. In Section II we review the established ${\cal O}\left(1/\sqrt{\omega}\right)$ anomaly for BD vacuum and Brans-Dicke-Maxwell electrovacuum. Section III introduces the Brans-Dicke-Klein-Gordon action, outlines the difficulties associated with the Jordan frame, and derives the field equations in the Einstein frame. Sections IV and V are dedicated to solving the Einstein-frame field equations in the harmonic radial coordinate. Section VI examines the solution at large $\omega$, while Section VII provides a formal proof of the observed ${\cal O}\left(1/\sqrt{\omega}\right)$ behavior. Section VIII concludes and offers an outlook on the implication of the ${\cal O}\left(1/\sqrt{\omega}\right)$ anomaly in the PPN formalism. The Appendix contains alternative presentations of the solution in various coordinate choices.
@@ II A recap of the ${\cal O}\left(1/\sqrt{\omega}\right)$ anomaly in Brans-Dicke vacuum
with $\bar{\phi}$ being a non-zero constant. This could be seen by expressing the BD scalar field and the metric components in Taylor series of $1/\omega$ as
@@
The $\mathcal{O}(1/\omega)$ behavior of the BD scalar field has faced challenges over the past thirty years. One concrete example central to this debate is the Brans Class I exact solution—a static spherisymmetric vacuum for BD gravity [32]. Previous investigations of this solution [14, 15, 16, 9, 10, 11, 12, 13, 25, 24, 22, 19, 18, 21, 20, 17, 23] have revealed that, in the limit of $\omega\rightarrow\infty$, (i) its metric configuration does not converge to a vacuum of General Relativity (GR), and (ii) its BD scalar field exhibits behavior in the form
@@
Moreover, it is important to note that, by virtue of the action (17), $\tilde{g}_{\mu\nu}$ is not a vacuum of GR. Rather, it is a configuration of GR sourced by a free massless scalar field (i.e., $\tilde{\phi}$). For the static spherisymmetric setup, $\tilde{g}_{\mu\nu}$ is known to be the FJNW spacetime [34, 37, 33, 35, 36] which deviates from the Schwarzschild spacetime.
@@
III Brans-Dicke action coupled with a Klein-Gordon field
@@
$\mathcal{S}$ $\simeq\sqrt{-g}\biggl{[}\phi\,\mathcal{R}-\frac{\omega}{\phi}g^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi-g^{\mu\nu}\partial_{\mu}\Psi\partial_{\nu}\Psi\biggr{]}$ (21)
@@ III.1 Difficulties in the Jordan frame
This subsection reveals the technical difficulties associated with the Jordan frame, and the need to move to the Einstein frame. To see this, let us consider a static spherisymmetric metric in the harmonic radial coordinate, viz.
@@
The relevant Ricci tensor components are
$\mathcal{R}_{tt}$ $=e^{4\gamma-2\alpha}\gamma^{\prime\prime}$ (28)
$\mathcal{R}_{uu}$ $=-\alpha^{\prime\prime}+\frac{1}{2}\alpha^{\prime 2}+\gamma^{\prime\prime}-2\gamma^{\prime 2}$ (29)
$\mathcal{R}_{\theta\theta}$ $=1+e^{-\alpha}\Bigl{(}\gamma^{\prime\prime}-\frac{1}{2}\alpha^{\prime\prime}\Bigr{)}$ (30)
@@
$\phi^{\prime\prime}(u)$ $=-\frac{(\Psi^{\prime}(u))^{2}}{2\omega+3}=-\frac{\Psi_{*}^{2}}{2\omega+3}$ (32)
$\Longrightarrow\ \ \phi(u)$ $=-\frac{\Psi_{*}^{2}}{2(2\omega+3)}u^{2}+Nu+v_{0}$ (33)
with $N$ and $v_{0}$ being integration constants. This simplicity is misleading, however. There is “a steep price to pay”: the operator $\nabla_{\mu}\nabla_{\nu}\phi$ in Eq. (22) typically involves the cumbersome Christoffel symbols. The nine non-zero Christoffel symbols are
${\displaystyle\begin{array}[]{lll}\Gamma_{tu}^{t}=\gamma^{\prime};\\ \Gamma_{tt}^{u}=\gamma^{\prime}e^{4\gamma-2\alpha};&\Gamma_{uu}^{u}=\alpha^{\prime}-\gamma^{\prime};\\ \Gamma_{\theta\theta}^{u}=(\gamma^{\prime}-\frac{1}{2}\alpha^{\prime})e^{-\alpha};&\Gamma_{\varphi\varphi}^{u}=\Gamma_{\theta\theta}^{u}\sin^{2}\theta;\\ \Gamma_{u\theta}^{\theta}=-\gamma^{\prime}+\frac{1}{2}\alpha^{\prime};&\Gamma_{\varphi\varphi}^{\theta}=-\cos\theta\sin\theta;\\ \Gamma_{u\varphi}^{\varphi}=-\gamma^{\prime}+\frac{1}{2}\alpha^{\prime};&\Gamma_{\theta\varphi}^{\varphi}=\frac{\cos\theta}{\sin\theta}\end{array}}$ (34)
The field equations for the metric components boil down to
$\left\{\begin{array}[]{rl}\mathcal{R}_{tt}&=\frac{1}{\phi}\nabla_{t}\nabla_{t}\phi-\frac{1}{\phi}\frac{g_{tt}(\nabla\Psi)^{2}}{2(2\omega+3)}\\ &=-\frac{1}{\phi}\Gamma_{tt}^{u}\phi^{\prime}-\frac{1}{\phi}\frac{g_{tt}g^{uu}\Psi^{\prime 2}}{2(2\omega+3)}\\ \mathcal{R}_{uu}&=\frac{\omega}{\phi^{2}}\phi^{\prime 2}+\frac{1}{\phi}\nabla_{u}\nabla_{u}\phi+\frac{1}{\phi}\left[\Psi^{\prime 2}-\frac{g_{uu}(\nabla\Psi)^{2}}{2(2\omega+3)}\right]\\ &=\frac{\omega}{\phi^{2}}\phi^{\prime 2}+\frac{1}{\phi}\left(\phi^{\prime\prime}-\Gamma_{uu}^{u}\phi^{\prime}\right)+\frac{1}{\phi}\frac{4\omega+5}{2(2\omega+3)}\Psi^{\prime 2}\\ \mathcal{R}_{\theta\theta}&=\frac{1}{\phi}\nabla_{\theta}\nabla_{\theta}\phi-\frac{1}{\phi}g_{\theta\theta}\frac{(\nabla\Psi)^{2}}{2(2\omega+3)}\\ &=-\frac{1}{\phi}\Gamma_{\theta\theta}^{u}\phi^{\prime}-\frac{1}{\phi}\frac{g_{\theta\theta}g^{uu}\Psi^{\prime 2}}{2(2\omega+3)}\end{array}\right.$ (35)
which subsequently lead to
@@
III.2 Transitioning from Jordan frame to Einstein frame
For the sake of illustration, instead of (21), we shall consider a BD action minimally coupled with a massless KG field $\Psi$ and a Maxwell EM field
$\sqrt{-g}\Bigl{[}\phi\,\mathcal{R}-\frac{\omega}{\phi}g^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi-g^{\mu\nu}\partial_{\mu}\Psi\partial_{\nu}\Psi-F_{\mu\nu}F^{\mu\nu}\Bigr{]}$ (37)
where the Maxwell EM field tensor is $F_{\mu\nu}:=\nabla_{\mu}A_{\nu}-\nabla_{\nu}A_{\mu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$. Let us define a new field $\theta$ per
@@ III.2 Transitioning from Jordan frame to Einstein frame
Then
$\sqrt{-g}\phi\mathcal{R}=\sqrt{-\tilde{g}}\Bigl{[}\tilde{\mathcal{R}}+3\tilde{g}^{\mu\nu}\tilde{\nabla}_{\mu}(\partial_{\nu}\theta)-\frac{3}{2}\tilde{g}^{\mu\nu}\partial_{\mu}\theta\partial_{\nu}\theta\Bigr{]}$ (41)
The middle term in right hand side of this last equation is
$\sqrt{-\tilde{g}}\tilde{g}^{\mu\nu}\tilde{\nabla}_{\mu}(\partial_{\nu}\theta)$ $=\sqrt{-\tilde{g}}\tilde{\nabla}_{\mu}\left(\tilde{g}^{\mu\nu}\partial_{\nu}\theta\right)$ (42)
$=\partial_{\mu}\left(\sqrt{-g}\tilde{g}^{\mu\nu}\partial_{\nu}\theta\right)$ (43)
which is a total derivative (by making use of the identity $\nabla_{\mu}A^{\mu}=\frac{1}{\sqrt{-g}}\partial_{\mu}(\sqrt{-g}A^{\mu})$). Therefore, we arrive at
@@
$\sqrt{-g}g^{\mu\nu}g^{\alpha\beta}F_{\mu\alpha}F_{\nu\beta}$ $=\sqrt{-\tilde{g}}\tilde{g}^{\mu\nu}\tilde{g}^{\alpha\beta}F_{\mu\alpha}F_{\nu\beta}$ (47)
@@
$\sqrt{-\tilde{g}}\Bigl{[}\tilde{\mathcal{R}}-\left(\omega+\frac{3}{2}\right)\tilde{g}^{\mu\nu}\partial_{\mu}\theta\partial_{\nu}\theta-e^{-\theta}\tilde{g}^{\mu\nu}\partial_{\mu}\Psi\partial_{\nu}\Psi\,$
$-\,\tilde{g}^{\mu\nu}\tilde{g}^{\alpha\beta}F_{\mu\alpha}F_{\nu\beta}\Bigr{]}$ (48)
Crucially, the Maxwell EM sector remains untransformed; it remains decoupled from the BD scalar field in the Einstein frame. In contrast, the KG field now forms a coupling with the BD scalar field, where the latter takes on the role of a “dilaton”, represented by the term $e^{-\theta}$. This introduces a level of complexity for the KG field.
III.3 Field equations of Einstein-Klein-Gordon-dilation (EKGD) action
In the rest of this paper, we shall not concern with the Maxwell EM sector in action (48). Instead, we shall focus on the following Einstein-Klein-Gordon-dilaton (EKGD) action:
$\sqrt{-g}\left[\mathcal{R}-g^{\mu\nu}\left(\left(\omega+3/2\right)\partial_{\mu}\theta\partial_{\nu}\theta+e^{-\theta}\partial_{\mu}\Psi\partial_{\nu}\Psi\right)\right]$ (49)
Note that, for this section, we have temporarily suppressed the tildes in (49) as compared with (48), for the sake of clarity. We shall restore the tildes in the later sections. Let us find the field equations for the action (49).
Upon variation, standard textbook derivation yields
$\delta\left(\sqrt{-g}\mathcal{R}\right)=\sqrt{-g}\left(\frac{1}{2}g^{\mu\nu}\mathcal{R}-\mathcal{R}^{\mu\nu}\right)\delta g_{\mu\nu}+\text{total deriv.}$. (50)
For other terms, denoting the symmetric rank-2 tensor $A_{\mu\nu}:=\left(\omega+3/2\right)\partial_{\mu}\theta\partial_{\nu}\theta+e^{-\theta}\partial_{\mu}\Psi\partial_{\nu}\Psi$ and $A:=g^{\mu\nu}A_{\mu\nu}$ the trace of $A_{\mu\nu}$, we have
$\delta\left(\sqrt{-g}g^{\mu\nu}A_{\mu\nu}\right)$
$=\sqrt{-g}\biggl{(}\frac{1}{2}g^{\alpha\beta}\delta g_{\alpha\beta}A+\delta g^{\mu\nu}A_{\mu\nu}+g^{\mu\nu}\delta A_{\mu\nu}\biggr{)}$ (51)
$=\sqrt{-g}\biggl{(}\frac{1}{2}g^{\alpha\beta}\delta g_{\alpha\beta}A-g^{\mu\alpha}g^{\nu\beta}\delta g_{\alpha\beta}A_{\mu\nu}+g^{\mu\nu}\delta A_{\mu\nu}\biggr{)}$ (52)
$=\sqrt{-g}\left(\frac{1}{2}g^{\mu\nu}A-A^{\mu\nu}\right)\delta g_{\mu\nu}+\sqrt{-g}g^{\mu\nu}\delta A_{\mu\nu}$ (53)
Regarding the tensor $A_{\mu\nu}$:
The first term is
$\delta\left(\partial_{\mu}\theta\partial_{\nu}\theta\right)$
$=\nabla_{\mu}(\delta\theta)\partial_{\nu}\theta+\partial_{\mu}\theta\nabla_{\nu}(\delta\theta)$ (55)
$=\left[\nabla_{\mu}\left(\delta\theta\partial_{\nu}\theta\right)-\nabla_{\mu}(\partial_{\nu}\theta)\delta\theta\right]$
$+\left[\nabla_{\nu}\left(\delta\theta\partial_{\mu}\theta\right)-\nabla_{\nu}(\partial_{\mu}\theta)\delta\theta\right]$ (56)
$=\left[\nabla_{\mu}\left(\delta\theta\partial_{\nu}\theta\right)+\nabla_{\nu}\left(\delta\theta\partial_{\mu}\theta\right)\right]-2(\nabla_{\mu}\nabla_{\nu}\theta)\,\delta\theta$ (57)
This leads to
$\sqrt{-g}g^{\mu\nu}\delta\left(\partial_{\mu}\theta\partial_{\nu}\theta\right)$
$=2\sqrt{-g}\nabla_{\mu}\left(\delta\theta\nabla^{\mu}\theta\right)-2\sqrt{-g}\square\,\theta\delta\theta$ (58)
[…]
and
Similarly to Eq. (59)
$\sqrt{-g}g^{\mu\nu}e^{-\theta}\delta\left(\partial_{\mu}\Psi\partial_{\nu}\Psi\right)$
$=2e^{-\theta}\partial{}_{\mu}\left(\sqrt{-g}\delta\Psi\nabla^{\mu}\Psi\right)-2\sqrt{-g}e^{-\theta}\square\,\Psi\,\delta\Psi$ (61)
[…]
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ -2\sqrt{-g}e^{-\theta}\square\,\Psi\,\delta\Psi$ (62)
leading to
$\sqrt{-g}g^{\mu\nu}\delta\left(e^{-\theta}\partial_{\mu}\Psi\partial_{\nu}\Psi\right)$
$=\sqrt{-g}e^{-\theta}\left[2(\partial_{\mu}\theta\nabla^{\mu}\Psi-\square\,\Psi)\delta\Psi-(\nabla\Psi)^{2}\,\delta\theta\right]$ (63)
Combining all terms, the variation of the EKGD action (49) is
$\sqrt{-g}\biggl{[}\left(\frac{1}{2}g^{\mu\nu}\mathcal{R}-\mathcal{R}^{\mu\nu}\right)-\left(\frac{1}{2}g^{\mu\nu}A-A^{\mu\nu}\right)\biggr{]}\,\delta g_{\mu\nu}$
$+\Bigl{[}(2\omega+3)\sqrt{-g}\square\,\theta+\sqrt{-g}e^{-\theta}(\nabla\Psi)^{2}\Bigr{]}\,\delta\theta$ (64)
$+2\sqrt{-g}e^{-\theta}(-\partial_{\mu}\theta\nabla^{\mu}\Psi+\square\,\Psi)\,\delta\Psi$
Imposing the extremality on (64), regarding $\delta g_{\mu\nu}$, we obtain
$\mathcal{R}^{\mu\nu}-\frac{1}{2}g^{\mu\nu}\mathcal{R}$ $=A^{\mu\nu}-\frac{1}{2}g^{\mu\nu}A$ (65)
which, upon taking the trace, $\mathcal{R}=A$, results in $\mathcal{R}_{\mu\nu}=A_{\mu\nu}$, or
$\mathcal{R}_{\mu\nu}=\left(\omega+3/2\right)\partial_{\mu}\theta\partial_{\nu}\theta+e^{-\theta}\partial_{\mu}\Psi\partial_{\nu}\Psi$ (66)
Regarding $\delta\theta$, Eq. (64) yields
$(2\omega+3)\,\square\,\theta=-e^{-\theta}(\nabla\Psi)^{2}$ (67)
Regarding $\delta\Psi$, Eq. (64) yields
$\square\,\Psi=\nabla^{\mu}\Psi\nabla_{\mu}\theta$ (68)
The last two equations can also be neatly cast as
$(2\omega+3)\,e^{-\theta}\,\square\,\theta$ $=-(e^{-\theta}\nabla^{\mu}\Psi)\,(e^{-\theta}\nabla_{\mu}\Psi)$ (69)
$\nabla^{\mu}\left(e^{-\theta}\nabla_{\mu}\Psi\right)$ $=e^{-\theta}\left(-\nabla^{\mu}\theta\nabla_{\mu}\Psi+\square\Psi\right)=0$ (70)
When a non-constant Klein-Gordon (KG) field $\Psi$ is present, Eq. (67) requires that the ’dilaton’ field $\theta$ must also be non-constant. It is important to emphasize that the field $\Psi$ is “self-gravitating”, meaning that it must be determined jointly with $g_{\mu\nu}$ and $\theta$ through the derived field equations.
IV Solving the BDKG field equations
Hereafter, we shall restore the tildes in the field equations (66), (69) and (70) to emphasize that they are in the Einstein frame, viz.:
@@ IV Solving the BDKG field equations
Assuming stationarity and spherisymmetry, we shall work in the harmonic radial coordinate in the Einstein frame, viz.

### v3

@@ Abstract
We present an exact static spherically symmetric solution for the Brans-Dicke action sourced by a self-gravitating massless Klein-Gordon helicity–0 field, which notably possesses an energy-momentum tensor with non-vanishing trace. Through a Weyl mapping into the Einstein frame, the Brans-Dicke scalar field transforms into a “dilaton” that couples with the Klein-Gordon field. Despite this dilatonic coupling, the field equations of the resulting Einstein-Klein-Gordon-dilaton action are fully soluble using the harmonic radial coordinate. Remarkably, in the limit of $\omega\rightarrow\infty$, the Brans-Dicke scalar field exhibits an anomalous behavior of $\mathcal{O}(1/\sqrt{\omega})$ as opposed to the expected $\mathcal{O}\left(1/\omega\right)$. Consequently, the solution converges to a spacetime configuration of General Relativity sourced by the original Klein-Gordon field and a free scalar field, the latter being the ${\cal O}\left(1/\sqrt{\omega}\right)$ “remnant” of the Brans-Dicke scalar field. Although this $\mathcal{O}\left(1/\sqrt{\omega}\right)$ anomaly was previously identified for Brans-Dicke vacuum and Brans-Dicke-Maxwell electrovacuum, our findings demonstrate its generic applicability in Brans-Dicke gravity, irrespective of the trace of the energy-momentum tensor of the source. We also provide a formal mathematical proof to bolster this conclusion. Furthermore, we discuss the potential implications of the $\mathcal{O}\left(1/\sqrt{\omega}\right)$ anomaly in improving the relativistic corrections to Newtonian gravity beyond the weak-field parametrized post-Newtonian formalism.
@@ I Motivation
This scenario is precisely what happens in BD gravity, revealing a non-convergence behavior that has been extensively explored and documented since the 1990s. Specifically, various authors have established that the infinite $\omega$ limit fails to transition the BD vacuum to a vacuum of GR [9, 10, 11, 12, 13, 14, 15, 16, 18, 19, 22, 20, 17, 21, 23]. Instead, it approaches a GR solution sourced by a free massless scalar field; (note: for static spherically symmetric setup, the limiting vacuum is the Fisher-Janis-Newman-Winicour (FJNW) solution  [^1] The solution has been re-discovered several times, and is also known as the Fisher-Bergmann-Leipnik-Janis-Newman-Winicour-Buchdahl-Wyman (FBLJNWBW) solution [23].). Moreover, the convergence rate in the limit is determined to be ${\cal O}\left(1/\sqrt{\omega}\right)$ instead of ${\cal O}\left(1/\omega\right)$. That is to say, the BD scalar field behaves as
@@
with $\bar{\phi}$ being a non-zero constant and the sub-leading field $\stackrel{{\scriptstyle 1}}{{\varphi}}$ non-constant. We shall aptly call $\stackrel{{\scriptstyle 1}}{{\varphi}}$ a “remnant” BD field in the infinite $\omega$ limit.
These findings are two sides of the same coin. Due to the $\stackrel{{\scriptstyle 1}}{{\varphi}}$ associated with the $1/\sqrt{\omega}$ order, the “kinetic” term remains finite, viz.
$\omega\left(\nabla\phi/\phi\right)^{2}\simeq\bigl{(}\ \nabla\stackrel{{\scriptstyle 1}}{{\varphi}}\>\bigr{)}^{2}$ (3)
@@
Secondly, we shall reinforce this proof by examining a concrete case: the BD theory sourced by a massless Klein-Gordon helicity-0 field. For the static spherically symmetric setup, this theory proves to be soluble, and the $\mathcal{O}\left(1/\sqrt{\omega}\right)$ signature manifests in the final analytical solution. As a by-product, the exact solution obtained herein can also serve as a prototype for future studies involving the BD action coupled with trace-carrying matter fields.
@@
Our paper is organized as follows. In Section II we review the established ${\cal O}\left(1/\sqrt{\omega}\right)$ anomaly for BD vacuum and Brans-Dicke-Maxwell electrovacuum. Section III introduces the Brans-Dicke-Klein-Gordon (BDKG) action, solves the Einstein-frame field equations in the harmonic radial coordinate system, and examines the solution at large $\omega$. Section IV provides a formal proof of the observed ${\cal O}\left(1/\sqrt{\omega}\right)$ behavior for a generic matter source. Sections V and VI derive the Eddington-Robertson-Schiff parameters $\beta$ and $\gamma$, and discuss the physical implications of our solution. Section VII concludes and offers an outlook on the implication of the ${\cal O}\left(1/\sqrt{\omega}\right)$ anomaly in the PPN formalism. Appendix A derives the field equations of the BDKG action for the Einstein frame. Appendix B expresses the solution in the isotropic coordinate system.
@@ II A recap of the ${\cal O}\left(1/\sqrt{\omega}\right)$ anomaly in Brans-Dicke vacuum
with $\bar{\phi}$ being a non-zero constant. The behavior in (10) is consistent with the (seemingly natural) following formal Taylor expansion of the solution
@@
The $\mathcal{O}(1/\omega)$ behavior of the BD scalar field has faced challenges over the past thirty years. One concrete example central to this debate is the Brans Class I exact solution—a static spherically symmetric vacuum for BD gravity [32]. Previous investigations of this solution [14, 15, 16, 9, 10, 11, 12, 13, 25, 24, 22, 19, 18, 21, 20, 17, 23] have revealed that, in the limit of $\omega\rightarrow\infty$, (i) its metric configuration does not converge to a vacuum of General Relativity (GR), and (ii) its BD scalar field exhibits behavior in the form
@@
Moreover, it is important to note that, by virtue of the action (17), $\tilde{g}_{\mu\nu}$ is not a vacuum of GR. Rather, it is a configuration of GR sourced by a free massless scalar field (i.e., $\tilde{\phi}$). For the static spherically symmetric setup, $\tilde{g}_{\mu\nu}$ is known to be the FJNW spacetime [34, 37, 33, 35, 36] which deviates from the Schwarzschild spacetime.
@@
III Brans-Dicke-Klein-Gordon action: New exact solution and $\mathcal{O}\left(1/\sqrt{\omega}\right)$ anomaly
@@
$\sqrt{-g}\biggl{[}\phi\,\mathcal{R}-\frac{\omega}{\phi}g^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi-g^{\mu\nu}\partial_{\mu}\Psi\partial_{\nu}\Psi\biggr{]}$ (21)
@@ III.1 Difficulties in the Jordan frame
This subsection reveals the technical difficulties associated with the Jordan frame, and the need to move to the Einstein frame. To see this, let us consider a static spherically symmetric metric in the harmonic radial coordinate, viz.
@@
$\phi(u)=-\frac{\Psi_{*}^{2}}{2(2\omega+3)}u^{2}+Nu+v_{0}$ (29)
with $N$ and $v_{0}$ being integration constants. This simplicity is misleading, however. There is “a steep price to pay”: the operator $\nabla_{\mu}\nabla_{\nu}\phi$ in Eq. (22) typically involves cumbersome Christoffel symbols. After full calculation of the Christoffel symbols and the Ricci tensor, the field equations leads to
@@
III.2 Mapping from Jordan frame to Einstein frame
For the sake of illustration, instead of (21), we shall consider a BD action minimally coupled with a massless KG field $\Psi$
$\sqrt{-g}\Bigl{[}\phi\,\mathcal{R}-\frac{\omega}{\phi}g^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi-g^{\mu\nu}\partial_{\mu}\Psi\partial_{\nu}\Psi\Bigr{]}\,.$ (31)
Let us define a new field $\theta$ per
@@ III.2 Mapping from Jordan frame to Einstein frame
These expressions give (modulo a total derivative term in the first identity below)
@@
$\sqrt{-\tilde{g}}\Bigl{[}\tilde{\mathcal{R}}-\left(\omega+\frac{3}{2}\right)\tilde{g}^{\mu\nu}\partial_{\mu}\theta\partial_{\nu}\theta-e^{-\theta}\tilde{g}^{\mu\nu}\partial_{\mu}\Psi\partial_{\nu}\Psi\Bigr{]}$ (38)
Crucially, the KG field now forms a coupling with the BD scalar field, where the latter takes on the role of a “dilaton”, represented by the term $e^{-\theta}$. This introduces a level of complexity for the KG field.
The field equations for the action (38) in the Einstein frame are derived to be (see Appendix A)
@@
III.3 Solving for the scalar fields
Assuming stationarity and spherical symmetry, we shall work in the harmonic radial coordinate  [^6] A harmonic radial coordinates satisfies the identity $\sqrt{-g}g^{11}=1$. in the Einstein frame, viz.